OpenStudy (anonymous):

what is y +2x -4 = x2 in general form?

OpenStudy (jhannybean):

\[\large y+ 2x -4 = x^2 \] in general form you could solve it for y, so be solving for y you would have to isolate the y on one side and your x-values on the other.

OpenStudy (jhannybean):

To do this subtract -2x from both sides and add +4 to both sides.

OpenStudy (campbell_st):

the standard form of a parabola is \[y = ax^2 + bx + c\]